OpenStudy (anonymous):

Algebra Help Please!!!!! Four pounds of onions costs the same as 2 pounds of string beans. At the same time, 1 pound of string bean costs 3 times as much as a pound of potatoes, while 1 pound of onion costs 4 cents less than 2 pounds of potatoes. What is the total cost (without tax) of 1 pound of each of the vegetables. Algebra Help Please!!!!! Four pounds of onions costs the same as 2 pounds of string beans. At the same time, 1 pound of string bean costs 3 times as much as a pound of potatoes, while 1 pound of onion costs 4 cents less than 2 pounds of potatoes. What is the total cost (without tax) of 1 pound of each of the vegetables. @Mathematics

OpenStudy (anonymous):

Let's let o represent onions, b represent beans, and p represent potatoes. 4o=2b b=3p o=2p-4 Now we substitute. 2o=b 2o=3p o=1.5p 1.5p=2p-4 -0.5p=-4 p=8 b=3(8)=24 4o=2(24)=48 o=12 All in cents

OpenStudy (anonymous):

omg thnx.......i think i get it now........